a copper plate 3mm thick has a cross sectional area 0.5. one face is at 80 degrees. while the other is 250 degrees. determine the rate of heat transfer through the plate. K=400
1
Expert's answer
2020-05-25T11:05:17-0400
"q=kA\frac{\Delta T}{\Delta x}"

"q=(400)(0.5)\frac{250-80}{0.003}=1.1\cdot10^7\ W"